OpenStudy (anonymous):

In a circle centered at point O, the ratio of the area of sector AOB to the area of the circle is 3/5 . What is the approximate measure, in radians, of the central angle corresponding to arc AB? Round the answer to two decimal places.

OpenStudy (ivancsc1996):

If you think about it, the ratio of the areas of AOB and the whole circle is the same as the ratio between the central angle corresponding to arc AB (theta) and the angle of the whole circle (2pi).\[3/5=\theta /2 \pi \rightarrow \theta=6\pi/5=\frac{ 6 }{ 5 }\pi\]

OpenStudy (anonymous):

what is that in a decimal??

OpenStudy (anonymous):

@ivancsc1996

OpenStudy (ivancsc1996):

3.78